A quarterback completes 12 out of 20 passes in his first game, how many passes
would you expect him to complete in his second game if he only throws 15 passes?

Found 2 solutions by MRperkins, stanbon:
Answer by MRperkins(300)

Email me at justin.sheppard.tech@hotmail.com with the problem and I will reply with the work in a .pdf file. We can also set up a time to work on problems together online using a whiteboard that will allow you to work through the problem by yourself. All of this is free (volunteer).

Answer by stanbon(75887)

A quarterback completes 12 out of 20 passes in his first game, how many passes
would you expect him to complete in his second game if he only throws 15 passes?
--------------------
Proportion:
x/15 = 12/20
x = 15(12/20)
x = 15(3/5)
x = 9
==============
Cheers,
Stan H.